The probability of an event is 4/7. Find the odds against this event occuring.
Answer by stanbon(75887)

The probability of an event is 4/7. Find the odds against this event occuring.
---------------
P(event) = 4/7
P(no event) = 3/7
----
Odd against = P(no even)/P(event) = (3/7)/(4/7) = 3:4
====================================
Cheers,
Stan H.
=============